theyadoreshayyy:

Point E is located at (2, −3), and point F is located at (−2, −1). Find the y-value for the point that is 3 over 4 the distance from point E to point F. −2.5 −2 −1 −1.5

theyadoreshayyy:

@Honda

Honda:

I can't help you on this one. sorry

theyadoreshayyy:

its okay thank you anyways

duhhitxmelivv:

Find the Distance between the y-coordinates of Point E and F. E(2.-3) F(-2,-1) y1=-3, y2 = -1 Calculate 3/4 of the distance between y-coordinates of E and F 3/4 * 2 = 3/2 = 1.5 We need to find the distance from Point E . That's why we add 1.5 to the y-coordinate E Y=-3+1.5= -1.5